# Regret Analysis of Posterior Sampling-Based Expected Improvement for Bayesian Optimization

### v2

@@ 1 Introduction
Bayesian optimization (BO) (Shahriari et al., 2016) is a powerful tool for optimizing expensive-to-evaluate black-box functions. BO aims for optimization with fewer observations due to the expensive evaluation cost of the objective function. For this purpose, BO sequentially queries a candidate determined by maximizing an acquisition function (AF) based on a Bayesian model. A Gaussian process (GP) model (Rasmussen and Williams, 2005) is typically employed in BO. BO has been applied to various fields, such as materials informatics (Ueno et al., 2016), robotics (Berkenkamp et al., 2023), and hyperparameter tuning (Snoek et al., 2012).
@@
Other lines of study have tackled theoretical guarantees on regret (for example, Srinivas et al., 2010; Russo and Van Roy, 2014). Seminal work (Srinivas et al., 2010) showed the cumulative regret upper bound for the GP upper confidence bound (GP-UCB) algorithm. Russo and Van Roy (2014); Chowdhury and Gopalan (2017) derived the cumulative regret upper bounds for the GP-based Thompson sampling algorithm (GP-TS). Furthermore, Scarlett (2018) showed the algorithm-independent regret lower bound and a tighter regret upper bound under several conditions. Recently, Takeno et al. (2024) showed that the variant of GP-PI called GP-PI from the maximum of sample path (GP-PIMS) achieves a sublinear Bayesian cumulative regret (BCR) upper bound. GP-PIMS avoids the problem of selecting the reference value by replacing it with the maximum of the posterior sample path, inspired by GP-TS. Furthermore, a recent study (Iwazaki, 2025) showed a tighter high-probability regret upper bound for the usual GP-UCB.
@@
This paper analyzes the posterior sampling-based GP-EI algorithm inspired by GP-PIMS (Takeno et al., 2024), which is GP-EI with the reference value set by the maximum of the posterior sample path. We refer to this algorithm as GP-EI from the maximum of the sample path (GP-EIMS). We show that GP-EIMS achieves the sublinear BCR under the Bayesian setting, where the objective function follows the GP. Moreover, GP-EIMS does not require rescaling of the posterior variance. Finally, we demonstrate the effectiveness of GP-EIMS, which stems from avoiding posterior variance rescaling, through numerical experiments.
@@ 2 Preliminary
where ${\cal X}\subset\mathbb{R}^{d}$ is an input domain and $d$ is a dimension. We assume that only observations $y=f(\boldsymbol{x})+\epsilon$, where $\epsilon$ is a noise, can be obtained, but are costly. Thus, our goal is to optimize $f$ with fewer observations. For this goal, BO sequentially performs querying an input $\boldsymbol{x}_{t}$ and observing $y_{t}=f(\boldsymbol{x}_{t})+\epsilon_{t}$, where $t$ is an iteration. The AF chooses queried input based on the Bayesian model, which is updated by the collected dataset.
@@
Furthermore, for the case of continuous ${\cal X}$, we assume the following smoothness condition for GP:
@@
where $L_{j}=\sup_{\boldsymbol{x}\in{\cal X}}\left|\frac{\partial f}{\partial\boldsymbol{x}_{j}}\right|$ and $[d]=\{1,\dots,d\}$.
@@
For regret analysis, we will use the quantity called maximum information gain (MIG) (Srinivas et al., 2010; Vakili et al., 2021):
@@
$\gamma_{T}\coloneqq\max_{A}I(\boldsymbol{y}_{A};\boldsymbol{f}_{A})\text{ such that }|A|=T\text{ and }\boldsymbol{a}_{i}\in{\cal X}\text{ for all }i\in[T],$ (4)
@@
For frequently used kernels, MIG is known to be sublinear (Srinivas et al., 2010; Vakili et al., 2021), for example, $\gamma_{T}=O\bigl{(}(\log T)^{d+1}\bigr{)}$ for the SE kernels and $\gamma_{T}=O\bigl{(}T^{\frac{d}{2\nu+d}}(\log T)^{\frac{2\nu}{2\nu+d}}\bigr{)}$ for the Matérn-$\nu$ kernels.
@@
We evaluate the theoretical performance of BO methods by the BCR (Russo and Van Roy, 2014; Kandasamy et al., 2018; Takeno et al., 2023, 2024) defined as
@@
where the expectation is taken with all the randomness, that is, $f,(\epsilon_{i})_{i\in[T]}$, and the randomness of the algorithm. Our goal is to show the sublinearity of the BCR since it implies that the simple regret $\mathbb{E}\left[f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{T})\right]\leq{\rm BCR}_{T}/T\rightarrow 0$ as $T\rightarrow\infty$ (Proposition 8 in Russo and Van Roy, 2018), where $\hat{\boldsymbol{x}}_{T}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{T}(\boldsymbol{x})$ is a recommendation point at $T$-th iteration.
@@
=$=\begin{cases}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{\mu_{t-1}(\boldsymbol{x})-y_{t-1}^{\rm max}}{\sigma_{t-1}(\boldsymbol{x})}\right)&\text{ if }\sigma_{t-1}(\boldsymbol{x})>0,\\ \max\{\mu_{t-1}(\boldsymbol{x})-y_{t-1}^{\rm max},0\}&\text{ if }\sigma_{t-1}(\boldsymbol{x})=0,\end{cases}$ (7)
@@ 3 Literature Review for Regret Analysis of GP-EI
The assumption for the regret analysis in the BO literature is twofold: the Bayesian setting that $f$ follows GPs (Srinivas et al., 2010; Russo and Van Roy, 2014; Takeno et al., 2023, 2024) and the frequentist setting that $f$ belongs to the known reproducing kernel Hilbert space (RKHS) (Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Janz et al., 2020; Iwazaki and Takeno, 2025). Furthermore, the analyzed regret can be categorized as cumulative regret $\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$ and simple regret $f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{t})$ with some recommended input at $t$-th iteration, for example, $\hat{\boldsymbol{x}}_{t}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$. In addition, as we discussed above, we refer to its expected variants as BCR and Bayesian simple regret in the Bayesian setting. Although our main analysis concentrates on the BCR analysis in the Bayesian setting, we review the analyses of GP-EI for both regrets in both settings.
@@ 3.2 Regret Analysis with Rescaling Posterior Variance
where $\nu_{t}=\Theta\left(\gamma_{t}\right)$. Their analysis shows $O(\gamma_{T}\sqrt{T})$ [^1] Although Theorem 1 in (Wang and de Freitas, 2014) states $O(\gamma_{T}^{3/2}\sqrt{T})$, we believe that their proof suggests $O(\gamma_{T}\sqrt{T})$. cumulative regret upper bound, which is sublinear, for example, for the Gaussian kernel. However, since $\nu_{t}\gg 1$ in most cases, the above rescaling strengthens the exploration behavior, and the practical effectiveness can deteriorate. As a result, in the practical scenarios, we must tune $\nu_{t}$ as a hyperparameter. Note that although Wang and de Freitas (2014) further consider a case where hyperparameters of GPs are unknown, we focus on a case where the hyperparameters of GPs are specified beforehand.
@@ 3.3 Other Related Works
Wang et al. (2025) tackle a theoretical analysis of the usual EI with $y_{t}^{\rm max}$ without the posterior variance scaling. However, Wang et al. (2025) analyze a quantity $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$, which is not the established regret definition, and can be negative. Paritularly when the noise variance $\sigma^{2}$ is large compared with $f(\boldsymbol{x}^{*})$, then $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$ can diverge to $\infty$- ∞ rapidly by noise. Therefore, analyzing $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$ is not suggestive in general.
@@ 4.1 Algorithm
Note that the posterior variance $\sigma_{t-1}^{2}(\boldsymbol{x})$ is not rescaled. Finally, an algorithm of GP-EIMS is shown in Algorithm 1.
@@ 4.2 Regret Analysis
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 00 otherwise.
@@
Fix $\beta>0$ and $U\leq\max_{\boldsymbol{x}\in{\cal X}}\bigl{\{}\mu_{t-1}(\boldsymbol{x})+\beta^{1/2}\sigma_{t-1}(\boldsymbol{x})\bigr{\}}$. Then, the following inequality holds:
@@
However, since the AF is defined by $U$ instead of $\max_{\boldsymbol{x}\in{\cal X}}\mu_{t-1}(\boldsymbol{x})$ in (Wang and de Freitas, 2014), we must further obtain an upper bound of $-2\log\left(\tau(-\beta^{1/2})\right)$. To obtain the upper bound, we leverage the following useful result:
@@
Hence, we see that $-2\log\left(\tau(-\beta^{1/2})\right)\leq\sqrt{2\pi\beta}-2\log(\phi(\beta^{1/2}))=\sqrt{2\pi\beta}+\beta+\log(2\pi)$. Consequently, by combining the results, we obtain the desired result. ∎
@@ 4.2.1 BCR Upper Bound for Finite Input Domain
Assume the same condition as in Lemma 4.4. Then, by running GP-EIMS, the following inequality holds with probability at least $1-\delta$:
@@ 4.2.2 BCR Upper Bound for Continuous Input Domain
For continuous input domains, as with the prior works (for example, Srinivas et al., 2010; Kandasamy et al., 2018; Takeno et al., 2023, 2024), we consider a discretized input set ${\cal X}_{t}\subset{\cal X}$, which is a finite set with each dimension equally divided into $m_{t}\geq 1$. Thus, $|{\cal X}_{t}|=m_{t}^{d}$. Furthermore, we define the nearest point of $\boldsymbol{x}$ in ${\cal X}_{t}$ as $[\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$.
@@
Figure 1: Results of simple regret for synthetic functions generated from a GP defined by the SE kernel. The top and bottom rows show the result on the length scale of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. The left, centered, and right columns show the result on the noise standard deviation $\sigma=0.01,0.1$, and $1$, respectively.
Figure 2: Results of cumulative regret for synthetic functions generated from a GP defined by the SE kernel. The top and bottom rows show the result on the length scale of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. The left, centered, and right columns show the result on the noise standard deviation $\sigma=0.01,0.1$, and $1$, respectively.
@@ 5 Experiments
As baselines, we employed GP-UCB (Srinivas et al., 2010), improved randomized GP-UCB (IRGP-UCB) (Takeno et al., 2023), which uses the random variable $\zeta_{t}$ instead of $\beta_{t}$ in GP-UCB, GP-TS (Russo and Van Roy, 2014), max-value entropy search (MES) (Wang and Jegelka, 2017), GP-PIMS (Takeno et al., 2024), GP-EI (Mockus et al., 1978), and GP-EI using the maximum of the posterior mean (Wang and de Freitas, 2014). We denote the method of (Wang and de Freitas, 2014) as GP-EI-$\mu^{\rm max}$. We set the hyperparameters of GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ to the theoretically derived parameters for the BCR analyses. Therefore, we set $\beta_{t}=2\log(|{\cal X}|t^{2}/\sqrt{2\pi}+1)$ for GP-UCB and $\zeta_{t}=2\log(|{\cal X}|/2)+Z_{t}$ with $Z_{t}\sim{\rm Exp}(\lambda=1/2)$ for IRGP-UCB  (Takeno et al., 2023). Furthermore, we set $\nu_{t}=2\log(|{\cal X}|t^{2}/\sqrt{2\pi}+1)$ for GP-EI-$\mu^{\rm max}$ (see Theorem B.4). The number of MC samples in MES was set to 10. We employed random Fourier feature approximations (Rahimi and Recht, 2008) for the posterior sampling in GP-EIMS, GP-TS, MES, and GP-PIMS.
@@
Figures 1 and 2 show the simple regret and cumulative regret for the SE kernels. We measured optimization performance by the simple regret $f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$, and the cumulative regret $\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$. We report the mean and standard error over 16 random trials for the generation of the initial dataset, the synthetic functions, the observation noise, and the algorithm’s randomness. The BO methods whose theoretical analysis has not been shown are indicated by dashed lines.
@@
GP-UCB and GP-EI-$\mu^{\rm max}$ consistently inferior to other methods. This is because the theoretical values of $\beta_{t}$ and $\nu_{t}$ are too large in practice. Indeed, IRGP-UCB, which uses $\zeta_{t}=O(\log|{\cal X}|)$ instead of $\beta_{t}=O(\log(|{\cal X}|t))$, shows better performance than GP-UCB and GP-EI-$\mu^{\rm max}$. However, IRGP-UCB slightly deteriorates when $\ell=0.2$, particularly for the cumulative regret, since $\zeta_{t}$ is still large for the objective function with $\ell=0.2$ whose local optima are not so many. Note that if $\ell$ becomes small, then the objective function has large variability and more local optima.
@@
GP-TS shows superior performance when $\ell=0.2$. However, when $\ell=0.1$, GP-TS often deteriorates for both simple and cumulative regret. This tendency has been reported as over-exploration of GP-TS (Takeno et al., 2023, 2024; Shahriari et al., 2016).
@@
MES shows superior performance in terms of cumulative regret, although its regret analysis has not been shown, as discussed in (Takeno et al., 2022). However, the simple regret stagnates in the case of $\ell=0.2$ and $\sigma=0.01,0.1$. Therefore, we can interpret that MES is slightly too exploitative in these experiments. On the other hand, it is known that GP-PIMS can be viewed as MES using one MC sample (Takeno et al., 2024). Superior performance of MES in terms of cumulative regret may imply the existence of a regret guarantee of MES.
@@
GP-EI shows superior performance for the simple regret, although its regret analysis for $\sigma>0$ has not been shown. However, in the experiments for the cumulative regret, GP-EI clearly deteriorates, particularly for the case of $\ell=0.2$. Since the noisy best observation $y^{\rm max}_{t}$ tends to be large with larger noise $\epsilon$, we can expect that GP-EI tends to lean more towards exploration. Since uncertainty sampling that maximizes $\sigma_{t-1}(\boldsymbol{x})$ achieves the simple regret upper bound that converges to zero (Vakili et al., 2021), it is understandable that GP-EI shows superior performance for the simple regret [^2] Although Vakili et al. (2021) analyze in the frequentist setting, their analysis immediately suggests a simple regret upper bound in the Bayesian setting.. On the other hand, we conjecture that the deterioration of GP-EI for the cumulative regret is due to over-exploration. Furthermore, we consider that since the objective function with $\ell=0.1$ has many local optima, the aggressive exploration by larger $y^{\rm max}_{t}$ results in relatively better performance compared with that of $\ell=0.2$.

### v3

@@ 1 Introduction
Bayesian optimization (BO) (Shahriari et al., 2016) is a powerful tool for optimizing expensive-to-evaluate black-box functions. BO aims for optimization with fewer observations due to the expensive evaluation cost of the objective function. For this purpose, BO sequentially queries a candidate input determined by maximizing an acquisition function (AF) based on some Bayesian model. That is, the queried candidate at $t$-th iteration $\boldsymbol{x}_{t}$ is chosen as $\boldsymbol{x}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}}\alpha(\boldsymbol{x})$, where $\alpha(\boldsymbol{x})$ is an AF computed based on the Bayesian model. A Gaussian process (GP) model (Rasmussen and Williams, 2005) is typically employed in BO. BO has been applied to various fields, such as materials informatics (Ueno et al., 2016), robotics (Berkenkamp et al., 2023), and hyperparameter tuning (Snoek et al., 2012).
@@
Other lines of study have tackled obtaining theoretical regret bounds (for example, Srinivas et al., 2010; Russo and Van Roy, 2014). Seminal work (Srinivas et al., 2010) showed the cumulative regret upper bound for the GP upper confidence bound (GP-UCB) algorithm. Russo and Van Roy (2014); Chowdhury and Gopalan (2017) derived the cumulative regret upper bounds for the GP-based Thompson sampling algorithm (GP-TS). Furthermore, Scarlett (2018) showed the algorithm-independent regret lower bound and a tighter regret upper bound under several conditions. Recently, Takeno et al. (2024) showed that the variant of GP-PI called GP-PI from the maximum of sample path (GP-PIMS) achieves a sublinear Bayesian cumulative regret (BCR) upper bound. GP-PIMS avoids the problem of selecting the reference value by replacing it with the maximum of the posterior sample path, inspired by GP-TS. Furthermore, a recent study (Iwazaki, 2025) showed a tighter high-probability regret upper bound for the usual GP-UCB.
@@
This paper analyzes the posterior sampling-based GP-EI algorithm inspired by GP-PIMS (Takeno et al., 2024), which is GP-EI with the reference value set by the maximum of the posterior sample path. We refer to this algorithm as GP-EI from the maximum of the sample path (GP-EIMS). We show that GP-EIMS achieves the sublinear BCR under the Bayesian setting, where the objective function follows GPs. Moreover, GP-EIMS does not require rescaling of the posterior variance. Finally, we demonstrate the effectiveness of GP-EIMS, which stems from avoiding posterior variance rescaling, compared with other theoretically guaranteed EI-based methods through numerical experiments.
@@ 2 Preliminary
where ${\cal X}\subset\mathbb{R}^{d}$ is an input domain and $d$ is a dimension. We assume that only observations $y=f(\boldsymbol{x})+\epsilon$, where $\epsilon$ is noise, can be obtained, but are costly. Thus, our goal is to optimize $f$ with the fewest observations possible. For this goal, BO sequentially performs queries on an input point $\boldsymbol{x}_{t}$ and observes $y_{t}=f(\boldsymbol{x}_{t})+\epsilon_{t}$, where $t$ is an iteration. The queried input $\boldsymbol{x}_{t}$ is chosen as $\boldsymbol{x}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\alpha(\boldsymbol{x})$, where $\alpha(\boldsymbol{x})$ is an AF computed based on the Bayesian model updated by the collected dataset.
@@
Furthermore, for the case of continuous $\cal X$}, we assume the following smoothness condition for GPs:
@@
where $L_{j}=\sup_{\boldsymbol{x}\in{\cal X}}\left|\frac{\partial f}{\partial\boldsymbol{x}_{j}}\right|$ and $d]=\{1,\dots,d\}$.
@@
For regret analysis, we will use the quantity called maximum information gain (MIG) (Srinivas et al., 2010; Vakili et al., 2021b):
@@
[…]
@@
For frequently used kernels, MIG is known to be sublinear (Srinivas et al., 2010; Vakili et al., 2021b), for example, $\gamma_{T}=O\bigl((\log T)^{d+1}\bigr)$ for the SE kernels and $\gamma_{T}=O\bigl(T^{\frac{d}{2\nu+d}}(\log T)^{\frac{2\nu}{2\nu+d}}\bigr)$ for the Matérn-$\nu$ kernels.
@@
We evaluate the theoretical performance of BO methods by the BCR (Russo and Van Roy, 2014; Takeno et al., 2023, 2024) defined as
@@
where the expectation is taken with all the randomness, that is, $f,(\epsilon_{i})_{i\in[T]}$, and the randomness of the algorithm. For example, in GP-TS, GP-PIMS, and GP-EIMS, the randomness of the algorithm implies the posterior sampling. Our goal is to show the sublinearity of the BCR since it implies that the simple regret $\mathbb{E}\left[f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{T})\right]\leq{\rm BCR}_{T}/T\rightarrow 0$ as $T\rightarrow\infty$ (Proposition 8 in Russo and Van Roy, 2018), where $\hat{\boldsymbol{x}}_{T}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{T}(\boldsymbol{x})$ is a recommendation point at $T$-th iteration. The Bayesian regret has been analyzed in many existing studies (Russo and Van Roy, 2014, 2018; Kandasamy et al., 2018; Paria et al., 2020; Takeno et al., 2023, 2024). Although the Bayesian regret quantifies the average performance of the algorithm, the Bayesian regret upper bound immediately implies a high probability upper bound due to Markov’s inequality. For more details, refer (Sec. 3.1 in Russo and Van Roy, 2014).
@@
$=\begin{cases}\sigma_{t-1}(\boldsymbol{x})\tau\left(\frac{\mu_{t-1}(\boldsymbol{x})-y_{t-1}^{\rm max}}{\sigma_{t-1}(\boldsymbol{x})}\right)&\text{ if }\sigma_{t-1}(\boldsymbol{x})>0,\\ \max\{\mu_{t-1}(\boldsymbol{x})-y_{t-1}^{\rm max},0\}&\text{ if }\sigma_{t-1}(\boldsymbol{x})=0,\end{cases}$ (7)
@@ 3 Literature Review for Regret Analysis of GP-EI
The two most commonly used assumptions for the regret analysis in the BO literature are the Bayesian setting that $f$ follows GPs (Srinivas et al., 2010; Russo and Van Roy, 2014; Takeno et al., 2023, 2024) and the frequentist setting that $f$ belongs to the known reproducing kernel Hilbert space (RKHS) (Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Janz et al., 2020; Iwazaki and Takeno, 2025). Furthermore, the analyzed regret can be categorized as cumulative regret $\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$ and simple regret $f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{t})$ with some recommended input at $t$-th iteration, for example, $\hat{\boldsymbol{x}}_{t}\coloneqq\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$. In addition, as we discussed above, we refer to its expected variants as BCR and Bayesian simple regret in the Bayesian setting. Although our main analysis concentrates on the BCR analysis in the Bayesian setting, we review the analyses of GP-EI for both regret definitions in both settings.
@@ 3.2 Regret Analysis with Rescaling Posterior Variance
where $\nu_{t}=\Theta\left(\gamma_{t}\right)$. Their analysis shows $O(\gamma_{T}\sqrt{T})$ [^1] Although Theorem 1 in (Wang and de Freitas, 2014) states $O(\gamma_{T}^{3/2}\sqrt{T})$, we believe that their proof suggests $O(\gamma_{T}\sqrt{T})$. cumulative regret upper bound, which is sublinear, for example, for the SE kernel. However, since $\nu_{t}\gg 1$ in most cases, the above rescaling strengthens the exploration behavior, and the practical effectiveness can deteriorate. As a result, in the practical scenarios, we must tune $\nu_{t}$ as a hyperparameter. Note that although Wang and de Freitas (2014) further consider a case where hyperparameters of GPs are unknown, we focus on a case where the hyperparameters of GPs are specified beforehand.
@@ 3.3 Other Related Works
Wang et al. (2025) tackle a theoretical analysis of the usual EI with $y_{t}^{\rm max}$ without the posterior variance scaling. However, Wang et al. (2025) analyze a quantity $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$, which is not the established regret definition, and can be negative. Paritularly when the noise variance $\sigma^{2}$ is large compared with $f(\boldsymbol{x}^{*})$, then $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$ can diverge to -$\infty$ rapidly by noise. Therefore, analyzing $f(\boldsymbol{x}^{*})-y^{\rm max}_{T}$ is not advisable in general.
@@ 4.1 Algorithm
Note that the posterior variance $\sigma_{t-1}^{2}(\boldsymbol{x})$ is not rescaled. Finally, an algorithm of GP-EIMS is shown in Algorithm 1. The difference from the usual AF of GP-EI is only changing the reference value as $g^{*}_{t}$ from $y^{\rm max}_{t-1}$.
@@ 4.2 Regret Analysis
where $C_{1}\coloneqq 2/\log(1+\sigma^{-2})$ and the indicator function $\mathbbm{1}\{\eta_{t}\geq 0\}=1$ if $\eta_{t}\geq 0$, and 0 otherwise.
@@
Fix $\beta>0$ and $U\leq\max_{\boldsymbol{x}\in{\cal X}}\bigl\{\mu_{t-1}(\boldsymbol{x})+\beta^{1/2}\sigma_{t-1}(\boldsymbol{x})\bigr\}$. Suppose that $k(\boldsymbol{x},\boldsymbol{x})\leq 1$ for all $\boldsymbol{x}\in{\cal X}$. Then, the following inequality holds:
@@
However, since the AF is defined by $U$ instead of $\max_{\boldsymbol{x}\in{\cal X}}\mu_{t-1}(\boldsymbol{x})$ in (Wang and de Freitas, 2014), we must further obtain an upper bound of $2\log\left(\tau(-\beta^{1/2})\right)$. To obtain the upper bound, we leverage the following useful result:
@@
Hence, we see that $2\log\left(\tau(-\beta^{1/2})\right)\leq\sqrt{2\pi\beta}-2\log(\phi(\beta^{1/2}))=\sqrt{2\pi\beta}+\beta+\log(2\pi)$. Consequently, by combining the results, we obtain the desired result. ∎
@@ 4.2.1 BCR Upper Bound for Finite Input Domain
Assume the same condition as in Lemma 4.4. Suppose that $k(\boldsymbol{x},\boldsymbol{x})\leq 1$ for all $\boldsymbol{x}\in{\cal X}$. Then, by running GP-EIMS, the following inequality holds with probability at least $1-\delta$:
@@ 4.2.2 BCR Upper Bound for Continuous Input Domain
For continuous input domains, as with the prior works (for example, Srinivas et al., 2010; Kandasamy et al., 2018; Takeno et al., 2023, 2024), we consider a discretized input set ${\cal X}_{t}\subset{\cal X}$, which is a finite set with each dimension equally divided into $m_{t}\geq 1$. Thus, $|{\cal X}_{t}|=m_{t}^{d}$. Furthermore, we define the nearest point of $\boldsymbol{x}$ in ${\cal X}_{t}$ as $\boldsymbol{x}]_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}^{\prime}\in{\cal X}_{t}}\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{1}$.
@@
$\mathrm{Pr}_{t}\left(\forall\boldsymbol{x}\in{\cal X},f(\boldsymbol{x})\leq\mu_{t-1}([\boldsymbol{x}]_{t})+2\beta_{t}^{1/2}(\delta/2)\sigma_{t-1}([\boldsymbol{x}]_{t})\right)\geq 1-\delta,$ (29)
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 1: Results of simple regret for synthetic functions generated from a GP defined by the SE kernel. The top, middle, and bottom rows show the result for noise standard deviations $\sigma=0.01,0.1$, and $1$, respectively. The left and right columns show the result for length scales of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. Daggers in the legend indicate that the BO method was performed with theoretical settings, for example, regarding the hyperparameters.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Results of cumulative regret for synthetic functions generated from a GP defined by the SE kernel. The top, middle, and bottom rows show the result for noise standard deviations $\sigma=0.01,0.1$, and $1$, respectively. The left and right columns show the result for length scales of the kernel $\ell=0.1$ and $\ell=0.2$, respectively. Daggers in the legend indicate that the BO method was performed with theoretical settings, for example, regarding the hyperparameters.
@@ 5 Experiments
As baselines, we employed GP-UCB (Srinivas et al., 2010), improved randomized GP-UCB (IRGP-UCB) (Takeno et al., 2023), which uses the random variable $\zeta_{t}$ instead of $\beta_{t}$ in GP-UCB, GP-TS (Russo and Van Roy, 2014), max-value entropy search (MES) (Wang and Jegelka, 2017), joint entropy search (JES) (Hvarfner et al., 2022; Tu et al., 2022), GP-PIMS (Takeno et al., 2024), GP-EI (Mockus et al., 1978), E3(Berk et al., 2019), and GP-EI using the maximum of the posterior mean (Wang and de Freitas, 2014). We denote the method of (Wang and de Freitas, 2014) as GP-EI-$\mu^{\rm max}$. We set the hyperparameters of GP-UCB, IRGP-UCB, and GP-EI-$\mu^{\rm max}$ to the theoretically derived parameters for the BCR analyses. Therefore, we set $\beta_{t}=2\log(|{\cal X}|t^{2}/\sqrt{2\pi}+1)$ for GP-UCB and $\zeta_{t}=2\log(|{\cal X}|/2)+Z_{t}$ with $Z_{t}\sim{\rm Exp}(\lambda=1/2)$ for IRGP-UCB  (Takeno et al., 2023). Furthermore, we set $\nu_{t}=2\log(|{\cal X}|t^{2}/\sqrt{2\pi}+1)$ for GP-EI-$\mu^{\rm max}$ (see Theorem B.4). The number of MC samples in MES, JES, and E3I was set to 10. We employed random Fourier feature approximations (Rahimi and Recht, 2008) for the posterior sampling in GP-EIMS, GP-TS, MES, JES, E3I, and GP-PIMS.
@@
Figures 1 and 2 show the simple regret and cumulative regret for the SE kernels. We measured optimization performance by the simple regret $f(\boldsymbol{x}^{*})-f(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t}(\boldsymbol{x})$, and the cumulative regret $\sum_{t=1}^{T}f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$. We report the mean and standard error over 16 random trials for the generation of the initial dataset, the synthetic functions, the observation noise, and the algorithm’s randomness.
@@
GP-UCB and GP-EI-$\mu^{\rm max}$ are consistently inferior to other methods. This is because the theoretical values of $\beta_{t}$ and $\nu_{t}$ are too large in practice. Indeed, IRGP-UCB, which uses $\zeta_{t}=\Theta(\log|{\cal X}|)$ instead of $\beta_{t}=\Theta(\log(|{\cal X}|t))$, shows better performance than GP-UCB and GP-EI-$\mu^{\rm max}$. However, IRGP-UCB slightly deteriorates when $\ell=0.2$, particularly for the cumulative regret, since $\zeta_{t}$ is still large for the objective function with $\ell=0.2$ whose local optima are not so many. Note that if $\ell$ becomes small, then the objective function has large variability and more local optima.
@@
GP-TS and JES show superior performance when $\ell=0.2$. However, when $\ell=0.1$, GP-TS and JES often deteriorate for both simple and cumulative regret. This tendency has been reported to be caused by over-exploration (Takeno et al., 2023, 2024; Shahriari et al., 2016).
@@
MES and E3I show superior performance in terms of cumulative regret, although their regret upper bounds have not been shown, as discussed in (Takeno et al., 2022) and Sec. 3. However, their simple regrets stagnate in the case of $\ell=0.2$ and $\sigma=0.01,0.1$. Therefore, we can interpret that MES and E3I are slightly too exploitative in these experiments since showing the relatively low cumulative regret and large simple regret implies that the corresponding BO method repeatedly evaluates non-optimal points whose instantaneous regret is low but not zero. On the other hand, it is known that GP-PIMS can be viewed as MES using one MC sample (Takeno et al., 2024). Furthermore, GP-EIMS is the same as E3I using one MC sample. Showing whether MES and E3I have regret guarantees is important future work.
@@
GP-EI shows superior performance for the simple regret, although its regret analysis for $\sigma>0$ has not been shown. However, in the experiments for the cumulative regret, GP-EI clearly deteriorates compared with E3I, MES, JES, GP-PIMS, and GP-EIMS, particularly for the case of $\ell=0.2$. Since the noisy best observation $y^{\rm max}_{t}$ tends to be large with larger noise $\epsilon$, we can expect that GP-EI tends to lean more towards exploration. Since uncertainty sampling that maximizes $\sigma_{t-1}(\boldsymbol{x})$ achieves the simple regret upper bound that converges to zero (Vakili et al., 2021a), it is understandable that GP-EI shows superior performance for the simple regret [^2] Although Vakili et al. (2021b) analyze in the frequentist setting, their analysis immediately suggests a simple regret upper bound in the Bayesian setting.. On the other hand, we conjecture that the deterioration of GP-EI for the cumulative regret is due to over-exploration. Furthermore, we consider that since the objective function with $\ell=0.1$ has many local optima, the aggressive exploration by larger $y^{\rm max}_{t}$ results in relatively better performance compared with that of $\ell=0.2$.